# Rui Lia,  Anyao Wangb   Mingqing Zhaic

### v2

@@ 4 Extremal graphs
([21]) Let $G$ be a connected graph of order $n$. Assume that $S$ is a non-empty proper subset of $V(G)$ and $T$ is a set of isolated vertices in $G-S$ such that $e(S,T)=|S||T|$. Let $H_{1},H_{2}$ be two graphs with the same vertex set $T$, and let $G_{i}$ be the graph obtained from $G$ by embedding the edges of $H_{i}$ into $T$. When $\rho(G)$ is sufficiently large compared to $|T|$, we have the following conclusions:
  • (i)
    If $H_{1}\equiv H_{2}$, then $\rho(G_{1})=\rho(G_{2})$.
  • (ii)
    If $H_{1}\succ H_{2}$, then $\rho(G_{1})>\rho(G_{2})$.
  • (iii)
    If $H_{1}\prec H_{2}$, then $\rho(G_{1})<\rho(G_{2})$.
@@
where $d_{H}(v)$ is the degree of vertex $v$ in $H$, and $p\geq 1$ is an integer. Ismailescu and Stefanica showed the following result regarding the maximization of $F_{p}(H)$.
@@
Considering that $p=2$ and $n=m-1$, similarly to Lemma 4.2 one can get the following result. For ease of understanding and to maintain integrity, we present the full proof here.
@@
Consider a graph $H^{*}\in\mathbb{G}(n,m)$ maximizing $F_{2}(H)$ over all graphs with $n$ vertices and $m$ edges. Let $(d_{1},d_{2},...,d_{n})$ represent the degrees of vertices $v_{1},v_{2},...,v_{n}$ in $H^{*}$. We first aim to show that the vertex of maximal degree in $H^{*}$ connects to all non-isolated vertices.
Suppose there exists a vertex $v_{i}$ not connected to another vertex $v_{j}$ with $d_{i}\leq d_{j}$, but both are connected to a third vertex $v_{k}$. By replacing the edge $v_{i}v_{k}$ with $v_{j}v_{k}$, we form a new graph $H^{\prime}\in\mathbb{G}(n,m)$. Then we have
$F_{2}(H^{\prime})-F_{2}(H^{*})=(d_{j}+1)^{2}+(d_{i}^{2}-1)-d_{j}^{2}-d_{i}^{2}=2d_{j}-2d_{i}+2>0,$
contradicting the maximality of $H^{*}$. Thus, the vertex of maximal degree must connect to all other non-isolated vertices.
Assume $v_{n}$ has maximal degree $d_{n}=l$ and is adjacent to $l$ vertices $\{v_{1},v_{2},...,v_{l}\}$. Since $m=n-1$, if $l<n-1$, it implies that $d_{l+1}=\cdots=d_{n-1}=0$. Our goal is to prove that $l=n-1$, which would make $H^{*}$ a star $K_{1,m}$.
Let $1\leq d_{1}\leq d_{2}\leq\cdots\leq d_{l}\leq d_{n}=l$. Suppose for contradiction that $l<n-1$. Then the subgraph induced by $\{v_{1},v_{2},...,v_{l}\}$ contains $n-1-l$ edges, it implies that $d_{l}>1$ and $d_{l}-1\leq n-1-l$.
We construct a new graph $H^{\prime\prime}$ from $H^{*}$ by removing the $d_{l}-1$ edges corresponding to $v_{l}$ from the subgraph with vertices $v_{1},\ldots,v_{l}$ and replacing them with $d_{l}-1$ edges between $v_{n}$ and $d_{l}-1$ isolated vertices. Note that we have sufficiently many isolated vertices to do this construction, since $d_{l}-1\leq n-1-l$.
The vertex degrees of $H^{\prime\prime}$ have changed as follows: $d^{\prime}_{l}=1,d^{\prime}_{l+1}=\cdots=d^{\prime}_{l+d_{1}-1}=1,d^{\prime}_{n}=l+d_{1}-1$, and for $11\leq i\leq l-$ 1, $d^{\prime}_{i}=d_{i}-1$ for $d_{l}-1$ vertices and $d^{\prime}_{i}=d_{i}$ for the others. We recall that $d_{i}\leq l$ for $1\leq i\leq l$. By direct computation, we obtain
$F_{2}(H^{*})-F_{2}(H^{\prime\prime})=\sum_{i=1}^{l-1}(d_{i}^{2}-d_{i}^{\prime 2})-(d_{l}-1)(2l-1).$
If $d_{i}^{\prime}=d_{i}-1$, then $d_{i}^{2}-d_{i}^{\prime 2}=2d_{i}-1\leq 2l-1$. It results that
$\sum_{i=1}^{l-1}(d_{i}^{2}-d_{i}^{\prime 2})\leq(d_{l}-1)(2l-1).$
Thus, we obtain that $F_{2}(H^{*})-F_{2}(H^{\prime\prime})\leq 0$ and $F_{2}(H^{*})-F_{2}(H^{\prime\prime})=0$ if and only if, for all indices $11\leq i\leq l-$ 1 for which $d_{i}^{\prime}=d_{i}-1$, we have $d_{i}=l$. Recall that $d_{i}\leq d_{l}$. Thus $d_{l}=l$, and this implies that all the vertices $v_{1},\ldots,v_{l-1}$, are neighbors of $v_{l}$. By the above observation, it results that $d_{1}=d_{2}=\cdots=d_{l}=l.$
Therefore, the only scenario in which our delete-add construction (performed under the assumption $d_{l}>1$) would not result in a graph $H^{\prime\prime}$ with a strictly higher $F_{2}$ value than our initial graph $H^{*}$ is when the only nontrivial component of $H^{*}$ is a complete graph on $l+1$ vertices.
In this case, we can compare directly $H=K_{l+1}\cup(n-1-l)K_{1}$ with our candidate $H^{*}=K_{1,m}$. We note that $m=\binom{l+1}{2}$, and obtain
$F_{2}(H)=(l+1)l^{2}\quad\text{and}\quad F_{2}(H^{*})=\binom{l+1}{2}^{2}+\binom{l+1}{2}.$
This implies that
$F_{2}(H^{*})-F_{2}(H)=\frac{(l-2)(l-1)l(l+1)}{4}\geq 0.$
This difference is non-negative, with equality holding only when $l=1$ or $l=2$. For $l=1$, $H^{*}$ becomes $K_{1,1}=K_{2}$, and for $l=2$, $H^{*}$ can also be $K_{3}\cup K_{1}$ when $m=3$.
Therefore, under the conditions specified, $H^{*}$ is uniquely the star $K_{1,m}$, except for the special case when $m=3$ where $K_{3}\cup K_{1}$ is also an extremal graph. ∎
@@
We now formally provide the proof of Theorem 1.3. Let $G$ be a spectral extremal graph that satisfies the property $P_{t,r}$. By Claim 3.2, we know that $V(G)=V\cup U$, where $|V|=\lfloor t\rfloor$ and $e(V,U)=|V||U|$. Let $T$ be a subset of $U$ such that $e(T,U\setminus T)=0$ and $|T|$ is constant. Consider two graphs $H_{1}$ and $H_{2}$ with $V(H_{1})=V(H_{2})=T$ and $e(H_{1})=e(H_{2})$. Define $G_{i}$ as the graph obtained from $G$ by embedding the edges of $H_{i}$ into $T$ for $i\in\{1,2\}$. Based on Lemma 4.1, we have the following observation.
Fact 4.1.
@@
We begin by establishing the following key claim.
@@
Suppose, for the sake of contradiction, that there exist two vertices $v_{1},v_{2}\in V$ such that $v_{1}v_{2}\not\in E(G)$. Since $\rho(G+v_{1}v_{2})>\rho(G)$, $G$ contains a subgraph $H$ such that $H+v_{1}v_{2}$ no longer satisfies the property $P_{t,r}$. We call $H+v_{1}v_{2}$ a counterexample of $G+v_{1}v_{2}$. Without loss of generality, we may assume that $H+v_{1}v_{2}$ is a minimal counterexample in terms of the number of vertices.
@@
First, we will show that $L\cup N_{1}$ is finite. Consider the size of the induced subgraph $G[V\cup B\cup L\cup N_{1}]$:
$e(V\cup B\cup L\cup N_{1})=e(H)+e(C,V(H))+e(C)+e(V,L\cup N_{1})+e(B,L)+e(L)+e(N_{1}).$
By the property $P_{t,r}$, we have
$e(V\cup B\cup L\cup N_{1})\leq t(|V|+|L|+|N_{1}|)+r.$
Note that $e(V(H))=\lfloor t|V(H)|+r\rfloor$, $e(B,L)+e(L)\geq|L|$, $e(N_{1})\geq|N_{1}|$. Therefore, it follows that
$e(V\cup B\cup L\cup N_{1})\geq\lfloor t|V(H)|+r\rfloor+|L|+|N_{1}|+(|L|+|N_{1}|)\lfloor t\rfloor.$
Combining the above two inequalities, we obtain
$(\lfloor t\rfloor+1-t)(|L|+|N_{1}|)$ $\leq t|V(H)|+r-\lfloor t|V(H)|+r\rfloor+t|C|$
$\leq 1+t|C|\leq 1+t\lfloor t\rfloor,$
@@
Now, we partition $N_{0}$ into two parts: Let $N_{0}^{1}$ denote the vertex set of all components with cardinality at least $1/t_{0}$, let $N_{0}^{2}$ denote the vertex set of all remaining components, where $t_{0}=\lfloor t\rfloor+1-t$.
Using a similar argument as before, we can show that $N_{0}^{1}$ is also finite. Consider the graph $G[V\cup B\cup N_{0}^{1}]$, we have
$e(V\cup B\cup N_{0}^{1})=e(V(H))+e(C,V(H))+e(C)+e(N_{0}^{1},V)+e(N_{0}^{1}).$
Given that $e(N_{0}^{1})$ for a forest $G[N_{0}^{1}]$ is at most $|N_{0}^{1}|-1$, and using the property $P_{t,r}$, we have
$e(V\cup B\cup N_{0}^{1})$ $\leq\lfloor t|V(H)|+r\rfloor+e(C,V(H))+e(C)+\lfloor t\rfloor|N_{0}^{1}|+|N_{0}^{1}|-1$
$\leq t(|V(H)|+|C|+|N_{0}^{1}|)+r.$
Since $20\leq|C|\leq\lfloor t\rfloor-$ 2, it follows that
$t_{0}|N_{0}^{1}|\leq t|C|-e(C,V(H))-e(C)+2<t^{2}.$
If $C\neq\emptyset$, we consider the graph $G[V\cup B]$:
$e(V\cup B)=e(V(H))+e(C,A)+e(C)+e(C,B).$
Using the property $P_{t,r}$, we get
$e(V\cup B)$ $=\lfloor t|V(H)|+r\rfloor+e(C,A)+e(C)+|C||B|$
$\leq t(|V(H)|+|C|)+r.$
Since $21\leq|C|\leq\lfloor t\rfloor-$ 2, we obtain
$|C||B|\leq t|C|-e(C,A)-e(C)+1\leq t|C|+1,$
which implies that
$|B|\leq t+1.$
Let $G^{\prime}$ be formed from $G$ by removing all edges induced by $B\cup L\cup N_{1}\cup N_{0}^{1}$ in $G$ and adding the edge $v_{1}v_{2}$. Since each component in $N_{0}^{2}$ has cardinality less than $1/t_{0}$, it is straightforward to verify that $G^{\prime}$ satisfies the property $P_{t,r}$.
@@
According to the proof of Theorem 2.1, we have $x_{u}\leq\frac{\lfloor t\rfloor}{\rho-s}x_{u^{*}}$ for $u\in U$, and $x_{v}\geq\big{(}1-\frac{1}{50s^{2}}\big{)}x_{u^{*}}$ for $v\in V$. Let $|B\cup L\cup N_{1}\cup N_{0}^{1}|=c$. Note that $\rho\geq\sqrt{\lfloor t\rfloor(n-\lfloor t\rfloor)}$. Then, we can deduce that
$\rho(G^{\prime})-\rho(G)$ $\geq X^{T}(A(G^{\prime})-A(G))X$
$=2x_{v_{1}}x_{v_{2}}-2\sum_{u_{i},u_{j}\in U\setminus N_{0}^{2}}x_{u_{i}}x_{u_{j}}$
$\geq 2\big{(}1-\frac{1}{50s^{2}}\big{)}^{2}x_{u^{*}}^{2}-2\big{(}\frac{c\lfloor t\rfloor}{\rho-s}\big{)}^{2}x_{u^{*}}^{2}$
$=2\big{[}\big{(}1-\frac{1}{50s^{2}}\big{)}^{2}-\big{(}\frac{c\lfloor t\rfloor}{\rho-s}\big{)}^{2}\big{]}x_{u^{*}}^{2}$
$>0,$
which is a contradiction since $G$ was assumed to have the maximum spectral radius.
If $C=\emptyset$, then we construct a new graph $G^{\prime\prime}$ as follows:
  • Delete all edges induced by $L\cup N_{1}\cup N_{0}^{1}$.
  • Delete all edges between $B$ and $L$.
  • Delete one edge in $G[B]$.
  • Add the edge $v_{1}v_{2}$.
The construction of $G^{\prime\prime}$ ensures that it still satisfies the property $P_{t,r}$ while potentially increasing the spectral radius, leading to a similar contradiction. Consequently, the contradictions derived from both cases force us to conclude that $V$ must form a clique, as desired. ∎

### v3

@@ 4 Characterization of extremal graphs
([22]) Let $G_{0}$ be an $n$-vertex connected graph. Assume that $S$ is a non-empty proper subset of $V(G_{0})$ and $T$ is a set of isolated vertices in $G_{0}-S$ such that $e(S,T)=|S||T|$. For $i\in\{1,2\}$, let $H_{i}$ be a graph with vertex set $T$, and let $G_{i}$ be the graph obtained from $G_{0}$ by embedding the edges of $H_{i}$ into $T$. When $\rho(G_{0})$ is sufficiently large compared to $|T|$, we have the following conclusions:
(i) If $H_{1}\equiv H_{2}$, then $\rho(G_{1})=\rho(G_{2})$.
(ii) If $H_{1}\succ H_{2}$, then $\rho(G_{1})>\rho(G_{2})$.
@@
where $p\geq 1$ is an integer. Ismailescu and Stefanica showed the following result regarding the maximization of $F_{p}(H)$.
@@
Given that $p=2$ and $n=m-1$, and following a similar approach to Lemma 4.2, the result below can be derived.
@@
Given a graph $F$, we define $\eta(F)=e(F)+(\lfloor t\rfloor-t)|V(F)|$, and let $\mathcal{C}(F)$ be the set of all components of $F$. It is easy to see that $\eta(F)=\sum_{F_{i}\in\mathcal{C}(F)}\eta(F_{i})$. Furthermore, a component $F_{i}$ of $F$ is said to be positive if $\eta(F_{i})>0$. Let $\mathcal{C}^{+}(F)$ be the set of all positive components of $F$. We denote
$\eta^{+}(F)=\sum_{F_{i}\in\mathcal{C}^{+}(F)}\eta(F_{i}),\ t_{0}=\lfloor t\rfloor+1-t\ \text{and}\ g_{0}=t\lfloor t\rfloor+r-\binom{\lfloor t\rfloor}{2}.$
Specially, if $\mathcal{C}^{+}(F)=\emptyset$, then $\eta^{+}(F)=0$. Recall that $r\geq-\binom{\lfloor t\rfloor+1}{2}$. Clearly, $t_{0}$ and $g_{0}$ are two constants with $0<t_{0}\leq 1$ and $g_{0}\geq 0$. Based on the definitions of $g_{0}$ and $\eta(F)$, we have the following two lemmas.
Lemma 4.4.
If $F^{\prime}$ is a subgraph of $F$, then $\eta(F^{\prime})\leq\eta(F)$.
@@
We first assume that $F$ is connected. If $V(F^{\prime})=V(F)$, then it is clear that $\eta(F^{\prime})\leq\eta(F)$. Next, suppose that $V(F^{\prime})\subsetneq V(F)$, and let $F_{1}^{\prime},\dots,F_{p}^{\prime}$ be the components obtained by removing $V(F^{\prime})$ from $F$. Since each $F_{i}^{\prime}$ is connected, $e(F_{i}^{\prime})\geq|V(F_{i}^{\prime})|-1$, and there is at least one edge connecting $F_{i}^{\prime}$ to $F^{\prime}$. Thus
$\eta(F)$ $=e(F)+(\lfloor t\rfloor-t)|V(F)|$
$\geq e(F^{\prime})+\sum_{i=1}^{p}e(F_{i}^{\prime})+\sum_{i=1}^{p}e(F_{i}^{\prime},F^{\prime})+(\lfloor t\rfloor-t)|V(F^{\prime})|+(\lfloor t\rfloor-t)\sum_{i=1}^{p}|V(F_{i}^{\prime})|.$
Combining $e(F_{i}^{\prime})\geq|V(F_{i}^{\prime})|-1$ and $e(F_{i}^{\prime},F^{\prime})\geq 1$, we have
$\eta(F)\geq\eta(F^{\prime})+\sum_{i=1}^{p}(|V(F_{i}^{\prime})|-1)+k+(\lfloor t\rfloor-t)\sum_{i=1}^{p}|V(F_{i}^{\prime})|=\eta(F^{\prime})+t_{0}\sum_{i=1}^{p}|V(F_{i}^{\prime})|,$
where $t_{0}=\lfloor t\rfloor+1-t>0$. It follows that $\eta(F^{\prime})\leq\eta(F)$.
Now we consider the case when $F$ is not connected. Let $\mathcal{C}(F)=\{F_{1},\dots,F_{q}\}$. Based on the above argument, we know that $\eta\big{(}F^{\prime}[V(F^{\prime})\cap V(F_{i})]\big{)}\leq\eta(F_{i})$ for $i\in\{1,\ldots,q\}$. It follows that $\eta(F^{\prime})\leq\sum_{i=1}^{q}\eta(F_{i})=\eta(F)$, as desired. ∎
Lemma 4.5.
$K_{\lfloor t\rfloor}\nabla F$ satisfies the property $P_{t,r}$ if and only if $\eta^{+}(F)\leq g_{0}$.
Proof.
First, assume that $K_{\lfloor t\rfloor}\nabla F$ satisfies the property $P_{t,r}$. Recall that $g_{0}=t\lfloor t\rfloor+r-\binom{\lfloor t\rfloor}{2}\geq 0$. If $\mathcal{C}^{+}(F)=\emptyset$, then $\eta^{+}(F)=0\leq g_{0}$, and the conclusion holds trivially. Otherwise, let $F^{+}$ be the union of all positive components of $F$. Since $K_{\lfloor t\rfloor}\nabla F^{+}$ is an induced subgraph of $K_{\lfloor t\rfloor}\nabla F$ with $|V(K_{\lfloor t\rfloor}\nabla F^{+})|\geq t$, it must satisfy $P_{t,r}$, and hence,
$e(K_{\lfloor t\rfloor}\nabla F^{+})=e(F^{+})+\lfloor t\rfloor|V(F^{+})|+\binom{\lfloor t\rfloor}{2}\leq t\big{(}\lfloor t\rfloor+|V(F^{+})|\big{)}+r.$
Rearranging this inequality, we obtain that $e(F^{+})+(\lfloor t\rfloor-t)|V(F^{+})|\leq g_{0},$ which implies that $\eta^{+}(F)=\eta(F^{+})\leq g_{0}$.
Now, assume that $\eta^{+}(F)\leq g_{0}$. We will prove that $K_{\lfloor t\rfloor}\nabla F$ satisfies the property $P_{t,r}$. Let $H$ be an arbitrary subgraph of $K_{\lfloor t\rfloor}\nabla F$ with $|V(H)|\geq t$. It suffices to show that $e(H)\leq t|V(H)|+r.$ Denote $F^{\prime}=F[V(H)\cap V(F)]$ and $t^{\prime}=|V(H)\cap V(K_{\lfloor t\rfloor})|$, where $0\leq t^{\prime}\leq\lfloor t\rfloor$. By Lemma 4.4, we have $\eta(F^{\prime})\leq\eta(F)\leq\eta^{+}(F)\leq g_{0},$ and hence
$e(F^{\prime})=\eta(F^{\prime})-(\lfloor t\rfloor-t)|V(F^{\prime})|\leq g_{0}-(\lfloor t\rfloor-t)(|V(H)|-t^{\prime}).$ (14)
Observe that $e(H)\leq e(F^{\prime})+\binom{t^{\prime}}{2}+t^{\prime}(|V(H)|-t^{\prime})$, and recall that $g_{0}=t\lfloor t\rfloor+r-\binom{\lfloor t\rfloor}{2}$. By combining (14), we obtain that
$e(H)\leq t|V(H)|+r-(\lfloor t\rfloor-t^{\prime})(|V(H)|-t^{\prime}-t)+\binom{t^{\prime}}{2}-\binom{\lfloor t\rfloor}{2}.$
Note that $\lfloor t\rfloor\geq t^{\prime}$ and $|V(H)|\geq t$. Define $f(t^{\prime})=(\lfloor t\rfloor-t^{\prime})t^{\prime}+\binom{t^{\prime}}{2}-\binom{\lfloor t\rfloor}{2}$. Then
$e(H)\leq t|V(H)|+r+f(t^{\prime}).$
If $t^{\prime}=\lfloor t\rfloor$, then $f(t^{\prime})=0$ and thus $e(H)\leq t|V(H)|+r$, as required. If $t^{\prime}\leq\lfloor t\rfloor-1$, it is easy to check that $f(\lfloor t\rfloor-1)=0$ and $f^{\prime}(t^{\prime})=\lfloor t\rfloor-t^{\prime}-\frac{1}{2}>0$. Thus $f(t^{\prime})\leq 0$ for $10\leq t^{\prime}\leq\lfloor t\rfloor-$ 1. Hence, $e(H)\leq t|V(H)|+r$. This completes the proof. ∎
@@
Let $G$ be a graph that attains the maximum spectral radius over all $n$-vertex graphs satisfying the property $P_{t,r}$. By Claim 3.2, we know that $V(G)=V\cup U$, where $|V|=\lfloor t\rfloor$ and $e(V,U)=|V||U|$. Let $T$ be a subset of $U$ such that $e(T,U\setminus T)=0$ and $|T|$ is constant. Define two new graphs $H_{1}$ and $H_{2}$ with $V(H_{1})=V(H_{2})=T$ and $e(H_{1})=e(H_{2})$. Let $G_{i}$ be the graph obtained from $G$ by embedding the edges of $H_{i}$ into $T$ for $i\in\{1,2\}$. By Lemma 4.1, we obtain the following corollary.
Corollary 4.1.
@@
Let $g(V)=t\lfloor t\rfloor+r-e(V)$. It is obvious that $g_{0}\leq g(V)\leq t\lfloor t\rfloor+r$. Observe that $G[V]\nabla F$ is a spanning subgraph of $K_{\lfloor t\rfloor}\nabla F$. By a similar argument as in the first part of Lemma 4.5, we obtain the following statement.
Corollary 4.2.
If $G[V]\nabla F$ satisfies the property $P_{t,r}$, then $\eta^{+}(F)\leq g(V).$
We will now formally present the proof of Theorem 1.3, starting with the following key claim. The proof proceeds via a minimal counterexample.
@@
Suppose to the contrary that $G[V]$ has a non-edge $v_{1}v_{2}$. Clearly, $\rho(G+v_{1}v_{2})>\rho(G)$. Then $G$ admits a subgraph $H$ of order at least $t$ such that $H+v_{1}v_{2}$ no longer satisfies the property $P_{t,r}$. Since $H$ satisfies the property $P_{t,r}$ but $H+v_{1}v_{2}$ no longer does, we conclude that $e(H)=\big{\lfloor}t|V(H)|+r\big{\rfloor}$. Now, $H+v_{1}v_{2}$ is called a counterexample of $G+v_{1}v_{2}$. We may assume that $H+v_{1}v_{2}$ is minimal with respect to the order.
@@
First, we will prove that the number of vertices in $L\cup N_{1}$ is finite. Consider the size of the induced subgraph $G[V\cup B\cup L\cup N_{1}]$:
$e(V\cup B\cup L\cup N_{1})$ $=e(V\cup B)+e(L\cup N_{1})+e(V\cup B,L\cup N_{1})$
$\geq e(H)+e(L)+e(N_{1})+e(B,L)+e(V,L\cup N_{1}).$
Recall that $e(H)=\big{\lfloor}t|V(H)|+r\big{\rfloor}$. Moreover, we see that $e(L)+e(B,L)\geq|L|$, $e(N_{1})\geq|N_{1}|$ and $e(V,L\cup N_{1})=\lfloor t\rfloor(|L|+|N_{1}|)$. Thus
$e(V\cup B\cup L\cup N_{1})\geq\big{\lfloor}t|V(H)|+r\big{\rfloor}+|L|+|N_{1}|+\big{\lfloor}t\big{\rfloor}\big{(}|L|+|N_{1}|\big{)}.$
On the other hand, by the property $P_{t,r}$, we have
$e(V\cup B\cup L\cup N_{1})\leq t(|V|+|B|+|L|+|N_{1}|)+r.$
Note that $|V|+|B|=|V(H)|+|C|$. Combining the above two inequalities, we obtain
$\big{(}\lfloor t\rfloor+1-t\big{)}\big{(}|L|+|N_{1}|\big{)}$ $\leq t\big{(}|V(H)|+|C|\big{)}+r-\big{\lfloor}t|V(H)|+r\big{\rfloor}$
$\leq t|C|+1\leq t\lfloor t\rfloor+1,$
@@
Recall that $t_{0}=\lfloor t\rfloor+1-t$. We now partition $N_{0}$ into two parts: Let $N_{0}^{1}$ denote the vertex set of all components in $G[N_{0}]$ with cardinality at least $\lfloor 1/{t_{0}}\rfloor+1$, and let $N_{0}^{2}$ denote the vertex set of all remaining components in $G[N_{0}]$.
Next, we will show that $|N_{0}^{1}|$ is also finite. Let $F$ be a component of $G[N_{0}^{1}]$. Note that $F$ is a tree with $|V(F)|\geq\lfloor 1/{t_{0}}\rfloor+1>1/{t_{0}}$, and recall that $\eta(F)=e(F)+(\lfloor t\rfloor-t)|V(F)|$. It follows that $\eta(F)=t_{0}|V(F)|-1>0$, and thus, $\eta^{+}(N_{0}^{1})=\eta(N_{0}^{1})=t_{0}|N_{0}^{1}|-c$, where $c$ is the number of the components in $G[N_{0}^{1}]$. Notice that $c(\lfloor 1/{t_{0}}\rfloor+1)\leq|N_{0}^{1}|$. Let $t^{\prime}_{0}=1/(\lfloor 1/{t_{0}}\rfloor+1)$. Then $t^{\prime}_{0}<t_{0}$, $c\leq t^{\prime}_{0}|N_{0}^{1}|$ and $\eta^{+}(N_{0}^{1})\geq(t_{0}-t_{0}^{\prime})|N_{0}^{1}|$. Since $G[V]\nabla G[N_{0}^{1}]$ satisfies the property $P_{t,r}$, by Corollary 4.2, we get that $\eta^{+}(N_{0}^{1})\leq g(V)\leq t\lfloor t\rfloor+r$. By combining these two inequalities, the assertion follows.
Thirdly, we shall demonstrate that $|B|$ is finite provided that $C\neq\emptyset$. Consider the subgraph $G[V(H)\cup C]$. It is easy to see that $e(V(H)\cup C)\geq e(H)+|C||B|$, where $e(H)=\big{\lfloor}t|V(H)|+r\big{\rfloor}\geq t|V(H)|+r-1$. However, using the property $P_{t,r}$, we have $e(V(H)\cup C)\leq t(|V(H)|+|C|)+r.$ Hence, $|C||B|\leq t|C|+1,$ and so $|B|\leq t+1$.
Observe that $e(B,N_{1}\cup N_{0}^{1}\cup N_{0}^{2})=0$. In view of (9), we have $e(B,L)\leq|L|(s-1)$. Building on the previous argument, we further see that both $e(B,L)$ and $e(L\cup N_{1}\cup N_{0}^{1})$ are constants. We proceed with the proof by considering two cases.
The first case is when $e(B)$ is finite. Now, $e(U\setminus N_{0}^{2})$ is also finite. Let $G^{\prime}$ be the graph obtained from $G$ by removing all edges within $U\setminus N_{0}^{2}$ and adding the edge $v_{1}v_{2}$. We assert that $G^{\prime}$ satisfies the property $P_{t,r}$. To verify this, observe that $E(G^{\prime}[U])=E(G^{\prime}[N_{0}^{2}])$. Let $F^{\prime}$ be a component of $G^{\prime}[U]$. Then $F^{\prime}$ is a tree with $|V(F^{\prime})|\leq\lfloor\frac{1}{t_{0}}\rfloor$. Thus
$\eta(F^{\prime})=e(F^{\prime})+(\lfloor t\rfloor-t)|V(F^{\prime})|=t_{0}|V(F^{\prime})|-1\leq 0,$ (15)
which implies that $G^{\prime}[U]$ contains no any positive components. Hence, $\eta^{+}(G^{\prime}[U])=0\leq g_{0}$. By Lemma 4.5, $K_{\lfloor t\rfloor}\nabla G^{\prime}[U]$ satisfies the property $P_{t,r}$. Since $G^{\prime}$ is a spanning subgraph of $K_{\lfloor t\rfloor}\nabla G^{\prime}[U]$, it follows that $G^{\prime}$ also satisfies the property $P_{t,r}$.
@@
By Theorem 2.1, $x_{v}\geq\big{(}1-\frac{1}{50s^{2}}\big{)}x_{u^{*}}$ for $v\in V$. In view of (9), we have $x_{u}\leq\frac{\lfloor t\rfloor}{\rho-s+1}x_{u^{*}}$ for each $u\in U$. Now we see that
$\rho(G^{\prime})-\rho(G)\geq X^{T}\big{(}A(G^{\prime})-A(G)\big{)}X=2x_{v_{1}}x_{v_{2}}-\!\!\sum_{u_{i}u_{j}\in E(G[U\setminus N_{0}^{2}])}\!\!2x_{u_{i}}x_{u_{j}}.$
Recall that $\rho\geq\sqrt{\lfloor t\rfloor(n-\lfloor t\rfloor)}$ and $e(G[U\setminus N_{0}^{2}])$ is finite. Therefore, we can deduce that $\rho(G^{\prime})>\rho(G)$, which contradicts the maximality of $\rho(G)$.
It remains the case when $e(B)$ is sufficiently large. Now $B$ is an infinite set, and by the previous proof, we have $C=\emptyset$. Thus $V(H)=V\cup B$. We construct $G^{\prime\prime}$ as follows: (i) delete all edges in $G[L\cup N_{1}\cup N_{0}^{1}]$; (ii) delete all edges from $L$ to $B$; (iii) delete an edge in $G[B]$ and add the edge $v_{1}v_{2}$ to $G[V]$. Using a similar argument to the one for $G^{\prime}$, we have $\rho(G^{\prime\prime})>\rho(G)$, which implies that $G^{\prime\prime}$ does not satisfy the property $P_{t,r}$.
We conclude the proof by contradiction, demonstrating that $G^{\prime\prime}$ satisfies the property $P_{t,r}$. Observe that $e(G^{\prime\prime}[V\cup B])=e(G[V\cup B])$. Since $G$ satisfies the property $P_{t,r}$, we have $e(G^{\prime\prime}[V\cup B])=e(G[V\cup B])\leq t(|V|+|B|)+r$. Recall that $V(H)=V\cup B$ and $H+v_{1}v_{2}$ is a minimal counterexample. We see that $e(H^{\prime})\leq t|V(H^{\prime})|+r$ for any subgraph $H^{\prime}$ of $G^{\prime\prime}[V\cup B]$. Hence, $G^{\prime\prime}[V\cup B]$ satisfies the property $P_{t,r}$.
On the other hand, observe that $G^{\prime\prime}[U\setminus B]\cong G^{\prime}[U\setminus B]$, and every component $F^{\prime}$ of $G^{\prime\prime}[U\setminus B]$ is a tree with $|V(F^{\prime})|\leq\lfloor\frac{1}{t_{0}}\rfloor$. In view of (15), we know that $\eta(F^{\prime})\leq 0$ for each component $F^{\prime}$ of $G^{\prime\prime}[U\setminus B]$. It follows that $\eta(G^{\prime\prime}[U\setminus B])\leq 0$.
Let $H^{\prime\prime}$ be an arbitrary subgraph of $G^{\prime\prime}$. Moreover, let $H_{1}^{\prime\prime}=G^{\prime\prime}[V(H^{\prime\prime})\cap(V\cup B)]$ and $H_{2}^{\prime\prime}=G^{\prime\prime}[V(H^{\prime\prime})\cap(U\setminus B)]$. Since $G^{\prime\prime}[V\cup B]$ satisfies the property $P_{t,r}$, we have $e(H_{1}^{\prime\prime})\leq t|V(H_{1}^{\prime\prime})|+r$. By Lemma 4.4, $\eta(H_{2}^{\prime\prime})\leq\eta(G^{\prime\prime}[U\setminus B])\leq 0$, which yields that $e(H_{2}^{\prime\prime})+\lfloor t\rfloor|V(H_{2}^{\prime\prime})|\leq t|V(H_{2}^{\prime\prime})|.$ Note that there are no edges between $B$ and $U\setminus B$ in $G^{\prime\prime}$. Thus, $e(H^{\prime\prime})\leq e(H_{1}^{\prime\prime})+e(H_{2}^{\prime\prime})+\lfloor t\rfloor|V(H_{2}^{\prime\prime})|\leq t|V(H^{\prime\prime})|+r.$ Therefore, $G^{\prime\prime}$ satisfies the property $P_{t,r}$. This completes the proof. ∎